Could someone please help me with this problem?
Part A.. Solve A=P+Prt for P if r=0.07, t=3, A=4646.40


Part B.. Solve A=1/2h(a+b)for b if a=3.2, h=12, A=51.6
Thank you for your help!!!!!!

Answer by jayanthihemkos@yahoo.com(87)

Part A.. Solve A=P+Prt for P if r=0.07, t=3, A=4646.40

Pluggin the values of r=0.07, t=3, A=4646.40 in
==> A=P+Prt
==> 4646.40 = p + p*0.07*3
==> 4646.40 = p + 0.21p
==> 4646.40 = p(1+0.21)
==> 4646.40 = p(1.21)
==> 4646.40/1.21 = p(1.21)/1.21
==> p = 4646.40/1.21
==> p = 3840